OpenStudy (anonymous):

fan and medal help please The function below models the correlation between the number of hours a plant is kept in sunlight (x) and the height (y), in mm, to which it grows: y = 4 + 2x What does the y-intercept of this function represent? The original height of the plant was 4 mm. The original height of the plant was 2 mm. The height of the plant increases by 2 mm for every hour of sunlight it receives. The height of the plant increases by 4 mm for every hour of sunlight it receives.

OpenStudy (anonymous):

@Savane0621 and @lizzyrockefeller

OpenStudy (anonymous):

do y= mx+b

OpenStudy (wmj259):

If you have x= 1 hour, you would add the 4 mm that it "born" with. So what do you think if x=0, whats y?

OpenStudy (anonymous):

Plant stays the same height. If x=0 the plant doesn't grow, it's original length was y=2 (mm) so it stays at that height.

OpenStudy (anonymous):

I dont know im retarded ... can someone walk me through the equation or show me how to do it,,, please

OpenStudy (wmj259):

It has 4mm already.

OpenStudy (wmj259):

If you graph that function it'll be y=2x+4. Tell me what you y-intercept is. My gosh.

OpenStudy (wmj259):

Your original function is y=2x+4. Right?

OpenStudy (anonymous):

yes

OpenStudy (wmj259):

y=mx+b Y is the output (mm) m is the slope of the line x is what you insert b is a CONSTANT (if x =0 what is left? B is left) So if you have zero hours of sunlight, what happens to the 4mm that is a constant?

OpenStudy (wmj259):

;)

OpenStudy (wmj259):

Good day.